OpenStudy (diamondanon2):

How many three-digit numbers greater than 200 can be formed from the digits 1, 2, 6, 7, and 9, if the digits can be repeated?

OpenStudy (diamondanon2):

Answer choices are 64 100 125

OpenStudy (diamondanon2):

@ShadowLegendX

TheSmartOne (thesmartone):

digits can be repeated _ _ _ we're looking a 3 digit number the first digit has to be 2 or greater we only have 2, 6, 7, or 9 the other two spots can be 1, 2, 6, 7, or 9

OpenStudy (diamondanon2):

I know it goes n! ---- (n-r)!r! I just dont know how to substitute

TheSmartOne (thesmartone):

for this one, we can use the counting princple what are the # of possible numbers that could go in the first digit? in the 2nd digit? 3rd digit? and you multiply it

TheSmartOne (thesmartone):

not sure how to use that formula here... it's easier to use the counting princple :P

OpenStudy (diamondanon2):

oh so that would be 100

OpenStudy (diamondanon2):

because 5^3

TheSmartOne (thesmartone):

5^3 = 125 but the first digit can only have 4 because if use 1 in the first digit the number will be less than 200 so 4 * 5 * 5

OpenStudy (diamondanon2):

oh okay that makes sense thank you so much!

OpenStudy (diamondanon2):

but 4 * 5 = 20 and 20 * 5 = 100?

OpenStudy (diamondanon2):

OH I SEE LUL MY BAD

OpenStudy (diamondanon2):

I got the correct answer by doing basic math incorrectly sorry

TheSmartOne (thesmartone):

yeah :)